OSG Technology Area Meeting, 5 February 2018¶
Coordinates: Conference: 719-284-5267, PIN: 57363; https://www.uberconference.com/osgblin
Attending: BrianL, Carl, Derek, Edgar, Marian, Mat, TimC
- OSG All Hands registration is open
- TimT out of office, returning Feb 19
- This week: Carl
- Next week: BrianL (?)
- 28 (+5) open tickets
|# of tickets||Δ||State|
|3||-21||Ready for Testing|
|0||-5||Ready for Release|
OSG Software Team¶
- Triage/LCMAPS VOMS tickets
- Make sure to update the "Next Action Deadline" after you or GOC staff have updated the ticket
- After a site transitions verify that pilots are still authenticating with factory ops
- Doc Focus
- Move tickets that haven't been started out of the sprint (https://jira.opensciencegrid.org/secure/RapidBoard.jspa?rapidView=7&projectKey=SOFTWARE)
- Update the doc tracking spreadsheet (https://docs.google.com/spreadsheets/d/1b3_9WqjUVlszu_tM23ehaOPQVAKb5OSRvmtuLx0u8Go/edit?usp=sharing)
after you've reviewed a doc and changes have been merged in
- Doc Focus Frenzy III: 2018-03-01, 1PM Central
None this week
- display.opensciencegrid.org (Carl) - able to reproduce the problem and has a fix
- FIU (BrianL) - completed LCMAPS VOMS transition
- Nebraska (Derek) - low CPU utilization fixed by downgrade to HTCondor 8.6.x
- OSC (Matyas) - Gave Trey an RSV patch to test
- SBGrid (Edgar) - assisting with a 3.2 -> 3.4 upgrade. Looking to transition off of their VOMS Admin server
OSG Release Team¶
|1||+1||Ready for Testing|
|0||+0||Ready for Release|
- GlideinWMS frontend bug fix
None this week
OSG Investigations Team¶
- CMS WLCG cpu utilization debugging
- Site issue with Nebraska. Fixed with update to HTCondor.
- Condor Ticket #6426
- Not going to worry about current records. Consulted with WLCG accounting team.
- OSG Docker images are getting a bit big, working with user support team to audit their contents
- Multiple Cuda versions were installed in single image.
- Doc Focus this week!
- New PS service coming to GRACC
- Currently running. Still need to redirect to GRACC's ES.
- HTCondor xfer stats
- Update with domain metrics.
- Ligo is going to run a StashCache origin server.
- Another round of data integrity on GRACC
- First focus was "Unknown" records in site records.
- Opened Ticket https://ticket.grid.iu.edu/35943
- FZU updated stashcache cache.
BrianL will follow up with Scott T to update HTCondor to a version that supports xfer stats